You own the following portfolio of stocks. Calculate each stock’s portfolio weight.

Stock Number of Shares Price per Share

A 650 $15.82   

B 320 $11.09

C 400 $39.80

D 100 $7.60

A: 33.70%

B: 11.63%

C: 52.18%

D: 2.49%

-Please check the answers and show work typed out no excel or grid style please as I am on mobile.
Answer #1

We need to calculate the weight of individual stock under portfolio.

weight of stock = Value of stock held in portfolio/Total portfolio value.

Computation of value of stock and portfolio

Stock A = 650*15.82 = 10283

Stock B = 320*11.09 = 3548.8

Stock C = 400*39.8=15920

Stock D = 100*7.6=760

Total value of portfolio = 10283+3548.8+15920+760= 30511.8

Weight of stock A = 10283/30511.8 = 33.7%

Weight of stock B = 3548.8/30511.8 = 11.63%

Weight of stock C = 15920/30511.8 = 52.18%

weight of Stock D = 760/30511.8= 249%